# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v2

@@ 4.1 Overall Performance on TLUE
Most models perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary models, the GPT series underperforms, whereas open-source models like Deepseek-v3 exceed the baseline. For reasoning-optimized models, Deepseek-r1 surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
@@
The performance gap between proprietary and open-source models is minimal, with Deepseek-v3 and Gemini-1.5-Flash performing comparably to Claude-3.5-Sonnet. Notably, GPT-3.5-Turbo underperforms relative to Llama3.1-8B-Instruct, suggesting that pretraining data, adaptation, and optimization significantly impact Tibetan language performance.
@@
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, Llama3.1-405B-Instruct excels in Other, and Qwen2.5-72b-Instruct in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
Performance on Ti-SafetyBench.
@@
Most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet achieves the highest accuracy (58.5% in CAA), whereas GPT-4O and GPT-3.5-Turbo underperform, with the latter significantly below random.
@@
Among open-source models, Deepseek-v3 surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. Llama3.1-405B-Instruct approaches the baseline, while smaller models like Llama3.1-8B-Instruct and Qwen2.5-72B-Instruct perform poorly. Reasoning-optimized models such as Deepseek-r1 excel in CAA but struggle in DA, whereas O1-mini consistently underperforms.
@@
Model Eval Method Avg. OFF UB PH MH IA EM PP
Claude-3.5-Sonnet DA 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini-1.5-Flash 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
GPT-4O 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT-3.5-turbo 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
Deepseek-v3 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
Llama3.1-405B-Instruct 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
Llama3.1-70B-Instruct 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
Qwen2.5-72B-Instruct 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen2.5-32B-Instruct 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen2.5-7B-Instruct 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
Llama3.1-8B-Instruct 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
Deepseek-r1 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
O1-mini 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Claude-3.5-Sonnet CAA 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini-1.5-Flash 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
GPT-4O 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT-3.5-turbo 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
Deepseek-v3 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
Llama3.1-405B-Instruct 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
Llama3.1-70B-Instruct 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
Qwen2.5-72B-Instruct 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen2.5-32B-Instruct 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen2.5-7B-Instruct 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
Llama3.1-8B-Instruct 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
Deepseek-r1 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
O1-mini 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
Table 2: Accuracy results of different LLMs on the Ti-SafetyBench dataset using two evaluation methods: Concern All Answer (CAA) and Direct Answer (DA). The full form of the 7 categories is provided in Appendix Table 4.
4.2 High-Resource vs. Low-Resource Performance
@@
To examine the performance disparity between high-resource and low-resource languages, we compare model accuracy on CMMLU (Chinese) Li et al. (2023) and Ti-MMLU (Tibetan) using the Concern All Answer (CAA) evaluation. Figure 3(a) illustrates performance differences across domains, while Appendix Table 5 presents detailed accuracy comparisons.
Performance Degradation.
@@
All models experience substantial accuracy drops from CMMLU to Ti-MMLU. Qwen2.5-72B-Instruct declines from 84.70% $\rightarrow$ 16.50%, GPT-4 from 68.90% $\rightarrow$ 17.51%, and ChatGPT from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%). This underscores the difficulty of adapting LLMs to low-resource languages.
Domain Performance Shifts.
@@
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU, models excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. Notable shifts include:
  • Qwen2.5-72B-Instruct achieves the highest accuracy in "Other" (87.35%) on CMMLU but shifts to "Humanities" (17.88%) on Ti-MMLU.
  • Llama3.1-70B-Instruct transitions from "Other" (74.72%) in CMMLU to "Social Sciences" (26.31%) in Ti-MMLU.
  • GPT-4’s best-performing category changes from "Other" (73.16%) in CMMLU to "Social Sciences" (19.69%) in Ti-MMLU(GPT-4O).
These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
STEM as the Weakest Domain.
@@
STEM remains the most challenging domain in both benchmarks. It consistently ranks lowest in CMMLU and Ti-MMLU, except for Llama3.1-70B-Instruct, which performs worst in "China-specific" tasks. This indicates significant difficulties in mathematical and technical reasoning in Tibetan.
Ranking Inconsistencies.
@@
Model rankings in CMMLU do not consistently translate to Ti-MMLU. For instance, Llama3.1-70B-Instruct ranks below Qwen2.5-72B-Instruct in CMMLU but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
Implications for Low-Resource Adaptation.
@@ 4.3 Safety in Low-Resource Languages
(a) Performance degradation from CMMLU (Chinese) to Ti-MMLU (Tibetan).
(b) Performance degradation from SafetyBench (Chinese) to Ti-SafetyBench (Tibetan)
Figure 3: Comparison of model performance on Chinese Benchmarks and TLUE across different tasks.
@@
To evaluate model safety in low-resource languages, we compare performance on SafetyBench (Chinese) Zhang et al. (2023) and Ti-SafetyBench (Tibetan) using the Concern All Answer (CAA) evaluation. Figure 3(b) illustrates accuracy drops across safety categories, while Appendix Table 6 provides detailed performance breakdowns. These results highlight the challenges of maintaining safety alignment in Tibetan.
Performance Degradation in Tibetan.
@@
Models show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 achieves 89.2% on SafetyBench, but GPT-4O, despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
Variability in Safety Categories.
@@
Performance degradation varies across safety categories. GPT-4 excels in Physical Health (PH) and Ethics and Morality (EM) on SafetyBench, whereas GPT-4O, despite being more advanced, performs best in Illegal Activities (IA) on Ti-SafetyBench, highlighting the challenges of transferring safety alignment to Tibetan.
@@
Similarly, Qwen-chat-7B outperforms random in EM and IA on SafetyBench, but Qwen2.5-7B-Instruct retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
@@
GPT-3.5-Turbo falls well below the random baseline on Ti-SafetyBench, while GPT-4O and Qwen2.5-7B-Instruct exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
Challenges in Low-Resource Safety Alignment.
@@
The sharp decline in performance on Ti-SafetyBench underscores the challenge of maintaining safety alignment in Tibetan. While models perform well on SafetyBench (Chinese), most fall below or barely match the baseline on Ti-SafetyBench. Even top-performing models struggle with safety generalization, highlighting the need for targeted fine-tuning and enhanced multilingual adaptation in low-resource environments.
4.4 Reasoning vs. Chat Models
@@
We compare reasoning-optimized and chat models on TLUE, analyzing response behavior, accuracy, and conditional accuracy. The results are summarized in Appendix Table 7811, with a broader comparison provided on Ti-MMLU (Appendix Table 9) and Ti-SafetyBench (Appendix Table 10), covering overall accuracy across knowledge and safety domains.
Response Rate and Output Behavior.
@@
Reasoning models, compared to chat models, tend to evaluate all answer choices before selecting the correct one, resulting in higher response rates for CAA than DA (Figure4(a), 4(b),Appendix Table 78). This behavior stems from a step-by-step evaluation process. When evaluating reasoning models, intermediate reasoning steps are filtered, and only the final output is considered, as outlined in Appendix Algorithm 3 and 4.
Impact of Response Rate on Performance.
@@
Lower response rates notably affect reasoning models’ performance, highlighting the inherent difficulty in responding to Tibetan prompts. The gap between DA and CAA accuracy underscores this challenge, with DA tasks requiring direct answer generation (Figure4(a), 4(b), Appendix Table 78). Reasoning models, unlike chat models, struggle more with Tibetan prompts due to their emphasis on logical inference over language generation.
Generalization of Reasoning Models in Low-Resource Languages.
@@
As shown in Appendix Table 11, Deepseek-r1 consistently outperforms Deepseek-v3 in conditional accuracy. Additionally, O1-mini outperforms GPT-4O in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini underperforms compared to GPT-4O in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
(a) Response rate, accuracy, and conditional accuracy on Ti-MMLU.
(b) Response rate and accuracy on Ti-SafetyBench.
Figure 4: Response rate, accuracy, and conditional accuracy for reasoning-optimized and chat models on TLUE.
@@ 4.5 Model Scale and Low-Resource Performance
We investigate the effect of model scale on Tibetan language understanding and safety alignment using Ti-MMLU (Figure 5(a) , Appendix Table 12) and Ti-SafetyBench (Figure 5(b), Appendix Table 13).
Effect of Scale on Tibetan Understanding.
Larger models generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LLaMA-3.1-405B-Instruct slightly outperforms LLaMA-3.1-70B-Instruct, while Qwen2.5-32B-Instruct surpasses Qwen2.5-72B-Instruct. Notably, Qwen2.5-7B-Instruct achieves comparable or superior accuracy to Qwen2.5-72B-Instruct in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
Impact of Scale on Safety Alignment.
Larger models show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen2.5-72B-Instruct significantly outperforms Qwen2.5-32B-Instruct, indicating that scaling benefits safety alignment more than general understanding. In contrast, LLaMA-3.1-70B-Instruct shows only marginal improvement over LLaMA-3.1-8B-Instruct, suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
Challenges in Scaling for Low-Resource Languages.
While larger models generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller models, such as Qwen2.5-7B-Instruct, achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.

### v3

@@ 5.1 Performance on Ti-MMLU
As shown in Table 1(a), most LLMs perform below the random baseline (25%), underscoring the challenge of Tibetan language understanding. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy, surpassing the baseline by 10.6 percentage points in CAA. Among proprietary LLMs, the GPT Achiam et al. (2023); OpenAI (2024) series underperforms, whereas open-source models like DeepSeek-V3 DeepSeek-AI et al. (2024) exceed the baseline. For reasoning-optimized LLMs, DeepSeek-R1 DeepSeek-AI et al. (2025) surpasses random performance in CAA but underperforms in DA, highlighting Tibetan’s difficulty as a low-resource language.
@@
The performance gap between proprietary and open-source models is minimal, with DeepSeek-V3 DeepSeek-AI et al. (2024) and Gemini-1.5-Flash DeepMind (2024) performing comparably to Claude-3.5-Sonnet Anthropic (2024). Notably, GPT-3.5-Turbo Achiam et al. (2023); OpenAI (2024) underperforms relative to LlaMA-3.1-8B Dubey et al. (2024), suggesting that pretraining data, adaptation, and optimization significantly impact Tibetan language performance.
@@
STEM remains the most challenging category, whereas most models perform best in Social Sciences. However, LlaMA-3.1-405B Dubey et al. (2024) excels in Other, and Qwen-2.5-72b Team (2024) in Humanities, indicating that structured reasoning tasks in Tibetan pose challenges for LLMs, while general knowledge and socially contextualized tasks are relatively easier.
The detial data is released in Appendix Table 15 and Appendix Table 16.
5.2 Performance on Ti-SafetyBench
@@
As shown in Table 1(b), most LLMs perform below the random baseline (36.7%), indicating significant challenges in aligning models with safety principles in Tibetan. Claude-3.5-Sonnet Anthropic (2024) achieves the highest accuracy (58.5% in CAA), whereas GPT-4O Achiam et al. (2023) and GPT-3.5-Turbo Achiam et al. (2023) underperform, with the latter significantly below random.
@@
Among open-source models, DeepSeek-V3 DeepSeek-AI et al. (2024) surpasses the baseline in both DA and CAA, demonstrating strong safety alignment. LlaMA-3.1-405B Dubey et al. (2024) approaches the baseline, while smaller models like LlaMA-3.1-8B Dubey et al. (2024) and Qwen-2.5-72B Team (2024) perform poorly. Reasoning-optimized models such as DeepSeek-R1 DeepSeek-AI et al. (2025) excel in CAA but struggle in DA, whereas O1-mini OpenAI (2024) consistently underperforms.
@@
5.3 High-Resource vs. Low-Resource
@@
To examine the performance disparity between high-resource and low-resource languages, we compare model accuracy on CMMLU Li et al. (2023) and Ti-MMLU using the CAA evaluation. Figure 2 illustrates performance differences across domains, while Appendix Table 6 presents detailed accuracy comparisons.
5.3.1 Performance Degradation
@@
All LLMs experience substantial accuracy drops from CMMLU Li et al. (2023) to Ti-MMLU. Qwen-2.5-72B Team (2024) declines from 84.70% $\rightarrow$ 16.50%, GPT-4 Achiam et al. (2023) from 68.90% $\rightarrow$ 17.51%, and ChatGPT Achiam et al. (2023); OpenAI (2024) from 53.22% $\rightarrow$ 3.42%, falling below the random baseline (25%), underscoring the difficulty of adapting LLMs to low-resource languages.
5.3.2 Domain Performance Shifts
@@
While accuracy declines across all domains, the highest-performing categories differ between languages. In CMMLU Li et al. (2023), LLMs excel in "Other" or "China-specific" categories, whereas in Ti-MMLU, "Social Sciences" and "Humanities" are strongest. For instance, Qwen-2.5-72B Team (2024) achieves the highest accuracy in "Other" (87.35%) on CMMLU Li et al. (2023) but shifts to "Humanities" (17.88%) on Ti-MMLU. LlaMA-3.1-70B Dubey et al. (2024) transitions from "Other" (74.72%) in CMMLU Li et al. (2023) to "Social Sciences" (26.31%) in Ti-MMLU. GPT-4’s Achiam et al. (2023) best-performing category changes from "Other" (73.16%) in CMMLU Li et al. (2023) to "Social Sciences" (19.69%) in Ti-MMLU (GPT-4O Achiam et al. (2023)). These shifts suggest that models retain general knowledge better but struggle with structured reasoning in Tibetan.
5.3.3 STEM as the Weakest Domain
@@
STEM remains the most challenging domain in both benchmarks. It consistently ranks lowest in CMMLU Li et al. (2023) and Ti-MMLU, except for LlaMA-3.1-70B Dubey et al. (2024), which performs worst in "China-specific" tasks. This indicates significant difficulties in mathematical and technical reasoning in Tibetan.
5.3.4 Ranking Inconsistencies
@@
LLM rankings in CMMLU Li et al. (2023) do not consistently translate to Ti-MMLU. For instance, LlaMA-3.1-70B Dubey et al. (2024) ranks below Qwen-2.5-72B Team (2024) in CMMLU Li et al. (2023) but surpasses it in Ti-MMLU, indicating that strong performance in high-resource languages does not necessarily predict effectiveness in low-resource settings.
5.3.5 Implication: Low-Resource Adaptation
@@ 5.4 Safety in Low-Resource Languages
(a) Performance Comparison of Qwen-2.5-72B Team (2024) from CMMLU Li et al. (2023) to Ti-MMLU
(b) Performance Comparison of LlaMA-3.1-70B Dubey et al. (2024) from CMMLU Li et al. (2023) to Ti-MMLU
(c) Performance Comparison of GPT-4o Achiam et al. (2023) from CMMLU Li et al. (2023) to Ti-MMLU
(d) Performance Comparison of GPT-3.5-turbo Achiam et al. (2023) from CMMLU Li et al. (2023) to Ti-MMLU
Figure 2: LLMs’ Performance Degradation from CMMLU Li et al. (2023) to Ti-MMLU
(a) Performance Comparison of GPT-4o Achiam et al. (2023) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
(b) Performance Comparison of GPT-3.5-turbo Achiam et al. (2023) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
(c) Performance Comparison of Qwen-2.5 (Chat) Team (2024) from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
Figure 3: LLMs’ Performance Degradation from SafetyBench Zhang et al. (2023) to Ti-SafetyBench
@@
To evaluate LLMs’ safety in low-resource languages, we compare performance on SafetyBench Zhang et al. (2023) and Ti-SafetyBench using CAA. Figure 3 illustrates accuracy drops across safety categories, while Appendix Table 7 provides detailed performance breakdowns. These results highlight the challenges of maintaining safety alignment in Tibetan.
5.4.1 Performance Degradation in Tibetan
@@
LLMs show a significant decline in safety alignment on Ti-SafetyBench, with most falling below the random baseline (36.7%). GPT-4 Achiam et al. (2023) achieves 89.2% on SafetyBench Zhang et al. (2023), but GPT-4O Achiam et al. (2023), despite being a more advanced model, drops to 32.9% on Ti-SafetyBench, particularly in Ethical and Moral Reasoning (92.6% $\rightarrow$ 24.8%). GPT-3.5-Turbo Achiam et al. (2023) experiences a sharper decline (80.4% $\rightarrow$ 11.6%), while Qwen-chat-7B Team (2024) decreases from 77.4% to 30.2%. These findings suggest that LLMs struggle with safety alignment in low-resource languages, often underperforming compared to random selection.
5.4.2 Variability in Safety Categories
@@
Performance degradation varies across safety categories. GPT-4 Achiam et al. (2023) excels in PH and EM on SafetyBench Zhang et al. (2023), whereas GPT-4O Achiam et al. (2023), despite being more advanced, performs best in IA on Ti-SafetyBench, highlighting the challenges of transferring safety alignment to Tibetan.
@@
Similarly, Qwen-chat-7B Team (2024) outperforms random in EM and IA on SafetyBench Zhang et al. (2023), but Qwen-2.5-7B Team (2024) retains its best performance only in EM on Ti-SafetyBench. Despite updates, both models show significant degradation in Tibetan.
@@
GPT-3.5-Turbo Achiam et al. (2023) falls well below the random baseline on Ti-SafetyBench, while GPT-4O Achiam et al. (2023) and Qwen-2.5-7B Team (2024) exceed or match the baseline in some categories. These results suggest that factors such as enhanced pretraining data, multilingual adaptation, and fine-tuning may contribute to better transferability, but safety alignment in low-resource settings remains a significant challenge.
5.4.3 Low-Resource Safety Alignment
@@
The sharp decline in performance on Ti-SafetyBench underscores the challenge of maintaining safety alignment in Tibetan. While models perform well on SafetyBench Zhang et al. (2023), most fall below or barely match the baseline on Ti-SafetyBench. Even top-performing models struggle with safety generalization, highlighting the need for targeted fine-tuning and enhanced multilingual adaptation in low-resource environments.
5.5 Reasoning vs. Chat
@@
We compare reasoning-optimized and chat LLMs on TLUE, analyzing response behavior, accuracy, and conditional accuracy. The results are summarized in Appendix C Table 8, Appendix C Table 9 and Appendix C Table 12, with a broader comparison provided on Ti-MMLU (Appendix C Table 10) and Ti-SafetyBench (Appendix C Table 11), covering overall accuracy across knowledge and safety domains.
5.5.1 Response Rate and Output Behavior
@@
Reasoning LLMs, compared to chat LLMs, tend to evaluate all answer choices before selecting the correct one, resulting in higher response rates for CAA than DA (Appendix C Table 8 and Appendix C Table 9). This behavior stems from a step-by-step evaluation process. When evaluating reasoning models, intermediate reasoning steps are filtered, and only the final output is considered, as outlined in Appendix B Algorithm 3 and Appendix B Algorithm 4.
5.5.2 Impact: Response Rate on Performance
@@
Lower response rates notably affect reasoning models’ performance, highlighting the inherent difficulty in responding to Tibetan prompts. The gap between DA and CAA accuracy underscores this challenge, with DA tasks requiring direct answer generation (Appendix C Table 8 and Appendix C Table 9). Reasoning models, unlike chat models, struggle more with Tibetan prompts due to their emphasis on logical inference over language generation.
5.5.3 Generalization of Reasoning Models in Low-Resource Languages
@@
As shown in Appendix C Table 12, DeepSeek-R1 Ranathunga et al. (2023) consistently outperforms DeepSeek-V3 DeepSeek-AI et al. (2024) in conditional accuracy. Additionally, O1-mini OpenAI (2024) outperforms GPT-4O Achiam et al. (2023) in STEM tasks, suggesting that stronger reasoning capabilities enhance performance in specialized tasks. However, O1-mini OpenAI (2024) underperforms compared to GPT-4O Achiam et al. (2023) in other domains, indicating that while reasoning optimization improves performance in low-resource settings, it is highly dependent on model architecture and optimization strategies. These results highlight the importance of both reasoning capabilities and generalization in low-resource language adaptation.
@@ 5.6 Model Scale and Low-Resource Performance
We investigate the effect of model scale on Tibetan language understanding and safety alignment using Ti-MMLU (Figure 4(a) , Appendix C Table 13) and Ti-SafetyBench (Figure 4(b), Appendix C Table 14).
5.6.1 Effect: Scale on Tibetan Understanding
LLMs generally yield better accuracy on Ti-MMLU, though improvements are inconsistent. LlaMA-3.1-405B Dubey et al. (2024) slightly outperforms LlaMA-3.1-70B Dubey et al. (2024), while Qwen-2.5-32B Team (2024) surpasses Qwen-2.5-72B Team (2024). Notably, Qwen-2.5-7B Team (2024) achieves comparable or superior accuracy to Qwen-2.5-72B Team (2024) in several domains, suggesting that model scaling alone does not guarantee enhanced performance in low-resource languages.
5.6.2 Impact: Scale on Safety Alignment
LLMs show improved performance on Ti-SafetyBench, particularly in categories involving complex ethical or factual reasoning. However, Qwen-2.5-72B Team (2024) significantly outperforms Qwen-2.5-32B Team (2024), indicating that scaling benefits safety alignment more than general understanding. In contrast, LlaMA-3.1-70B Dubey et al. (2024) shows only marginal improvement over LlaMA-3.1-8B Dubey et al. (2024), suggesting that model architecture and adaptation strategies are crucial for safety alignment in low-resource settings.
5.6.3 Challenge: Scaling for Low-Resource
While LLMs generally perform better, their advantage is less pronounced in Tibetan compared to high-resource languages. Smaller LLMs, such as Qwen-2.5-7B Team (2024), achieve competitive results in specific tasks, indicating that effective pretraining and fine-tuning strategies can mitigate the limitations of smaller models in low-resource environments.